OpenStudy (anonymous):

write 10 3/7 as an improper fraction A. 30/7 B. 13/7 c. 73/7 d. 37/7

OpenStudy (imsleeziboii):

c

OpenStudy (imsleeziboii):

10x7=70+3=73/7

OpenStudy (imsleeziboii):

medall

OpenStudy (study312):